OpenStudy (anonymous):

How many different arrangements of the letters in the word BETTER are there? answers 180 360 720

OpenStudy (lgbasallote):

Better has 6 letters e repeats twice t repeats twice so.. \[\large \frac{6!}{2!2!} = \frac{6 \times 5 \times 4 \times 3 \times 2}{2\times 2} = \frac{6 \times 5 \times 4 \times 3}{2}\] can you take it from there?

OpenStudy (anonymous):

so you get 720

OpenStudy (lgbasallote):

not exactly...you can simplify it further into \[6 \times 5 \times 3 \times 2\]

OpenStudy (anonymous):

i mean 360

OpenStudy (lgbasallote):

360 is just 6 x 5 x 4 x 3 you have to divide by 2

OpenStudy (anonymous):

my bad be cause u divide it by 2

OpenStudy (lgbasallote):

so...the final answer is...

OpenStudy (anonymous):

no that equal 720

OpenStudy (anonymous):

because you have to do 6*5*4*3*2

OpenStudy (lgbasallote):

\[\large \frac{6!}{2!2!} = \frac{6 \times 5 \times 4 \times 3 \times 2}{2 \times 2}\]

OpenStudy (lgbasallote):

do you get it?

OpenStudy (anonymous):

h ok my bad on top you put 6! ----- 2

OpenStudy (lgbasallote):

on the bottom it's 4 becaus 2 x 2 is 4

OpenStudy (anonymous):

on the bottom not 2*2

OpenStudy (lgbasallote):

what do you mean?

OpenStudy (anonymous):

u put 2 on the bottom not 2*2

OpenStudy (lgbasallote):

oh..that was because i canceled one of the two in the numerator and denominator...so i got \[\large \frac{6 \times 5\times 4 \times 3}{2}\] see one of the two in the numerator and one of the 2 in the denominator is missing?

OpenStudy (anonymous):

ok well thank you